Light of 10 cm wavelength is classified as what type of radiation?

x-ray

microwave

UV

γ-ray

visible
Answer #1

Range of wavelength of radiation is given by,

X - Ray = 1*10^-11 m to 1*10^-8 m

UV = 1*10^-8 m to 4*10^-7 m

Y -ray = Gamma rays = 1*10^-15 m to 1*10^-11 m

Visible = 4*10^-7 m to 7*10^-7 m

Microwave = 1*10^-3 m to 1*10^-1 m

given, wavelength = 10 cm = 1*10^-1 m

So, this Light is in microwave region.

therefore correct option is B.)

Please upvote.